Dear Richard dear all,

please find below a short summary of the activities done Sunday & today on the mountain related to the AO system. Italian people working at LBTO were:

Sunday 28/02
Arcetri: A. Riccardi, M. Xompero, R. Briguglio
total # of person: 3

Monday 11/03
Arcetri: A. Puglisi, P. Stefanini, E. Pinna, A. Riccardi, M. Xompero, R. Briguglio

Total # of person: 6

W unit

Activity of 28/02

none

Activity of 01/03

1) test with AGW rotation to verify the position of the W alignment mirror done with Andrew and the laser tracker. The off-axis error of our alignment mirror has been estimated to be less then 1 arcmin. This test required to place the AGW unit on some stable wood blocks to imprve the accuracy of the measurement. The error estimated is below the required accuracy.
2) The WFS board cleaning before telescope installation has started. 80% of the parts has been cleaned. A problem has been encountered with a grain of dust spoiling the image of one of the pupils. The attempt to remove this dust grain has not been successful yet.


AO control SW

No activity

LBT672a

28/02
- started functionality tests
- the recalibration of FF matrix showed act#220 having large FF matrix fitting error because instability. The cause is under investigation.
- shell flattening command (pre-calibrated in Arcetri) applied: peak force of 0.20N (std spec) is obtained pre-filtering command pattern using first 400 modes
- control gain set to the values pre-calibrated in Arcetri: the mirror self-oscillate with astigmatic shape at about 620Hz and edge amplitude of 70nm rms
- the self-oscillation disappears lowering proportional gain from 0.13N/um to 0.10N/um.

01/03
- removed the connections to the coldplate of the poles for mounting the shell: the autoscillations reduced the amplitude of a factor 2 (with full gain 0.13N/um)
- mounting the miniskirt (telescope configuration) the autoscillation disappers.
- run of the modal step responce test: longest settling time 0.94ms, largest overshoot: 6.3%. All within specifications and same results as in Arcetri
- with the shell set and flattened we ran the elevation slew test, from z=0 to z= 80deg, no reset of the shell. When moving back to zenith the elevation mechanism caused a large oscillation of the unit and the safety diagnostic reset the shell. Data under reductions.

Logistic

Reaching the telescope today from the base camp required about 2.5 hours because of state road conditions. The Arcetri group come down at 4:30 in the afternoon.

Others

A meeting has been held in the morning of 01/03 to prepare a second version of the schedule after collecting some more informations. The schedule has been sent to J. Brynnel for revision. A final version will be released on Wednesday.

Next update tomorrow

All the best

Simone

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback